### v2

# GaSLight: Gaussian Splats for Spatially-Varying Lighting in HDR

Abstract
We present GaSLight, a method that generates spatially-varying lighting from regular images. Our method proposes using HDR Gaussian Splats as light source representation, marking the first time regular images can serve as light sources in a 3D renderer. Our two-stage process first enhances the dynamic range of images plausibly and accurately by leveraging the priors embedded in diffusion models. Next, we employ Gaussian Splats to model 3D lighting, achieving spatially variant lighting. Our approach yields state-of-the-art results on HDR estimations and their applications in illuminating virtual objects and scenes. To facilitate the benchmarking of images as light sources, we introduce a novel dataset of calibrated and unsaturated HDR to evaluate images as light sources. We assess our method using a combination of this novel dataset and an existing dataset from the literature. The code to reproduce our method will be available upon acceptance.
1 Introduction
Lighting is crucial not just for shaping an image, but also for ensuring realism and aesthetic appeal when inserting virtual content. Estimating accurate lighting representations from one or many photographs is vital for creating lifelike visuals that convincingly blend within their surroundings and convey realistic surface appearance. This is essential in numerous applications including multimedia productions, virtual reality, forensics, cultural displays, and more.
Despite its importance, current methods face significant limitations in capturing and representing lighting accurately. One reason for this is the difficulty in capturing high dynamic range (HDR), which require multiple exposures [10] or specialized apparatus [58, 41]. Most approaches [69, 36, 12] therefore attempt to convert regular, low-dynamic range (LDR) images to HDR. However, as will be demonstrated later, they primarily focus on extending the dynamic range for tonemapping and visualization purposes [62, 20, 24], rather than extrapolating the intensity of bright light sources.
Another reason is the difficulty of simultaneously estimating HDR lighting that is both spatially-varying and high frequency. Tab. 1 provides an overview of existing lighting representations that have been proposed in the literature. Existing approaches either focus on accurate energy estimation, using parametric light sources [17], which allow for near-field effects, such as positioning a discrete lamp in the scene. Alternatively, per-pixel spherical harmonics [18] or tend to struggle in modeling high-frequency lighting and cannot generate strong cast shadows. Volumetric representations [32, 57, 66, 33] do not accurately model high-frequency lighting variations, as needed for reflections on glossy surfaces and do not readily integrate within existing rendering engine. Another prevalent representation used by lighting estimation methods is the environment map [16, 56, 55, 9, 44] (commonly referred to as IBL, for image-based lighting), which models infinitely-distant lighting as a spherical image. This representation cannot yield near-field lighting effects, such as a lamp in the scene, and model lighting as an infinitely-distant sphere. These limitations reduce the versatility of current lighting estimation methods, impacting their ability to achieve realistic lighting in scenes.
Param. SH Vol. IBL Ours
Near-field
Spatially-varying
HF reflections
HF lighting
Easy
Schematic [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 1: Comparison of different lighting representations. We include parametric (point or directional light sources, “param.”), Spherical Harmonics (“SH”), volumetric (“Vol.”), environment maps (“IBL”), and our proposed HDR GS representation. Aspects compared are near-field effects, such as the ability to represent a light source within the scene, High-Frequency reflections allowing for high-resolution details in reflections, High-Frequency lighting allowing strong shadows, and its overall ease of use (e.g., integration in existing rendering engines).
In this work, we introduce GaSLight, a novel method that addresses these challenges by capturing the high-frequency and spatially-varying lighting of a scene. It allows for near-field lighting effects, and readily integrates within existing rendering engines. GaSLight employs HDR Gaussian Splats as light representation, which is reconstructed from a set of LDR images. As opposed to HDRGS [6] (or HDR-NeRF [27]), our approach does not require alternating exposures to recover the HDR—rather we rely on a novel, specially-designed LDR-to-HDR network based on powerful diffusion models. Compared to previous NeRF-based HDR estimation approaches [19], our method is more accurate, does not require per-scene finetuning, and is much faster.
In summary, our contributions are threefold. First, we present a framework for using Gaussian Splats as a spatially-varying HDR light representation that can be estimated from a few images of a scene. This light representation can readily be used in a 3D renderer, enabling accurate and versatile lighting effects. Second, at the heart of our framework is a novel LDR-to-HDR estimation model, which leverages the strong priors in diffusion models to plausibly estimate the intensity of light sources. Finally, we compare our method on existing HDR datasets from the literature, and show that our methods best predicts the intensity of bright light sources. We also demonstrate its applicability for virtual object insertion where virtual objects can be inserted at multiple locations in the scene and relit appropriately in a multi-view setting.
2 Related work
Inverse tonemaping
has been an active area of research for several decades, with early methods focusing on inverting the camera function through various ad-hoc extrapolation techniques [25, 1]. Later, this topic was revisited using deep learning, first using end-to-end approaches [12, 13, 37], then by explicitly inverting the camera pipeline [36]. Inverse tonemapping has also been applied to spherical 360° panoramas, both outdoors [72] and indoors [69].
More recently, generative imaging has been utilized in inverse tonemapping, with GlowGAN [60] being the first to recover a multi-exposure bracket, as in traditional HDR capture [10]. Exposure Diffusion [3] improves this idea by adding exposure constraints to a pretrained diffusion model to generate relative exposures. While it capitalizes on the robust priors found in diffusion models, this method is restricted to pixel-space diffusion. Subsequently, LeDiff [62] finetunes a latent diffusion model [50] to determine relative exposures and merge them into an image with an extended dynamic range. Although this method shows promising results with latent-space diffusion for inverse tonemapping, it remains constrained by the decoder’s restricted dynamic range, which does not fully represent the richness of natural lighting. Our method draws inspiration from recent diffusion-based HDR estimation to propose an iterative dynamic range expansion scheme, enabling a significantly greater increase in dynamic range than current approaches allow.
3D scene representations
The way 3D content is represented has evolved significantly over time, addressing different aspects such as geometry and appearance. Traditional methods often relied on textured 3D meshes [15] as the primary means to represent scene geometry and appearance. Implicit radiance functions, such as those introduced in Neural Radiance Fields (NeRFs) [38], have revolutionized 3D scene representations by leveraging neural networks to represent complex geometries and appearances, successfully modeling complex, view-dependent effects. HDR-NeRF [27], HDR-HexPlane [68], and others [28, 39, 65] have extended this framework to handle HDR content, allowing for the reconstruction of scenes with an extended dynamic range. PanoHDR-NeRF [19] combine an LDR-to-HDR network trained for 360° panoramas with an omnidirectional NeRF representation to capture the spatially-varying HDR light field of an indoor scene.
Later, explicit radiance functions based on 3D Gaussian Splatting [29] were proposed. HDR-GS [6] and CineGR [61], leveraged multi-exposure input photographs to capture an extended dynamic range, similarly to their predecessor HDR-NeRF. As opposed to these methods, we focus on the extrapolation of the bright light sources present in the scene and do so without the requirement of multi-exposure inputs. We therefore create an explicit GS-based HDR representation, allowing them to function effectively as light sources for 3D renderers.
Camera pose estimation
To capture scenes in the wild, a common first step is to estimate the relative camera pose between pictures taken of a given scene. The classical way to achieve this is through bundle adjustment [59], a method which has been revisited recently by COLMAP [52, 53]. Through time, classical feature descriptors used for matching made way to learned descriptors [11, 51], which evolved to end-to-end learned relative camera pose estimation methods such as Mickey [2], Dust3r [64] and its follow-up, Mast3r [31]. We encourage the reading of the survey provided in [63].
Lighting representations
have evolved significantly over time, each addressing different aspects of light modeling and scene illumination. Parametric lights, such as point lights or directional lights, have been foundational since the inception of computer graphics. Methods that learn to predict parametric lights from images have been proposed [17, 71, 22], providing intuitive control over lighting but often struggle with accurately representing complex environments.
Spherical harmonics offer an efficient representation for lighting in real-time applications. They have been used to model both irradiance [47, 21] and radiance [54]. Garon et al. [18] estimate spatially-varying spherical harmonics directly from a single image.
Environment maps [4] provide a non-parametric alternative for capturing complex lighting environments, as they represent an infinitely distant light source modeled as a spherical image. They have been employed in many lighting estimation methods [16, 56, 26, 55, 19, 9, 44]. Implicit volumetric lighting representations were also devised to model both indoor [57, 33] and outdoor [67] lights, or for inverse rendering purposes [34]. Their opaqueness thwarts their powerful modeling capabilities, making them challenging to interpret and edit. In contrast, using Gaussian splats as a lighting representation combines the strengths of both parametric and non-parametric approaches. By representing HDR Gaussians with light-emitting spherical harmonics, our method provides spatially-varying lighting capabilities while maintaining the flexibility and interpretability of parametric representations.
3 Method
Our method takes a set of regular LDR images as input and produces lighting as an HDR 3D Gaussian Splats (3DGS) scene. The 3DGS can then be queried to obtain the amount of incoming radiance from a specific direction for rendering purposes. We break down our method into two steps. First, we uplift the LDR image to unsaturated HDR, providing plausible intensities for the observed light sources. Then, we fit the 3DGS scene to the estimated HDR content. It is important to note that our method does not perform lighting estimation from unseen parts of the scene—rather, it models only the lights visible within at least one image. We first describe our per-image light intensity estimation method, followed by the 3D scene reconstruction we employ.
3.1 HDR light intensity estimation
Refer to caption Refer to caption
(a) training (b) inference
Figure 1: Overview of our HDR estimation pipeline. (a) We train a diffusion model to predict two exposures (brighter or darker) given an input image. (b) At inference time, we recursively apply the same network to obtain an exposure stack, which is subsequently merged to HDR.
Our HDR estimation method aims to produce a plausible hallucination of the intensities of the bright light sources present in an LDR image. As illustrated in fig. 1, our approach relies on a diffusion model $\mathcal{D}$, which accepts as input an image at a certain exposure $I_{0}$, and is trained (fig. 1a) to predict two versions of the input image at different exposures: $I_{-2}=\mathcal{D}_{-}(I_{0})$, and $I_{+2}=\mathcal{D}_{+}(I_{0})$. Here, the $I_{X}$ notation means applying a $2^{X}$ factor to the exposure of the input image $I_{0}$. The resulting exposures (fig. 1b) are then merged to a single HDR image using a conventional HDR merging algorithm [10].
Multi-exposure prediction network
[…]
At inference time, the network can be applied recursively to the predicted exposures to further expand the dynamic range (i.e., $I_{+4}=\mathcal{D}_{+}(\mathcal{D}_{+}(I_{0}))$ as illustrated in fig. 1b. In practice, we repeat this process until images no longer contain (under- or over-)saturated pixels.
HDR merging
We consider the output from our multi-exposure estimation as a sequence of exposure-bracketed images $\{I_{-2n},...,I_{+2m}\}$ [48], where $n$ and $m$ indicate the number of dark and bright exposures estimated by the network, respectively. We first linearize the images ($\gamma=2.2)$, and then apply the Debevec HDR fusion method [10] to combine the sequence into a single HDR image. We adjust the weighting function to give a weight of $1$ to pixels with intensity values in the range of $[0.2,0.8]$ in the original image $I_{0}$, ensuring unsaturated pixels are preserved as much as possible from $I_{0}$. This step prevents blurring due to the averaging of slightly different pixel values that can be generated by the diffusion model across exposures. The rest of the pixels are weighted using the original weighting function, $w\!=\!-2\left\lvert z-0.5\right\rvert\!+\!1$. This results in an image in linear, unbounded colorspace.
360° prediction
Our method can also be naively applied to 360° panoramas by using a cube map representation and applying our method on each of the six faces independently. While this does not explicitly ensure consistency for light sources crossing the edges of the cube’s face, we observe that results obtained this way are usable and produce plausible lighting and reflections, effectively converting 360° into HDRI maps.
Training data and implementation details
We employ a dataset of 12,611 HDR 360° panoramas to train our light intensity estimation model, containing a mix of publicly available [45, 5] and captured data. At training time, we project the spherical panorama to an image plane, using random parameters drawn according to the distributions shown in tab. 2. Note that images are first re-exposed using [49] such that the of the log-luma channel is 0.17 prior to applying the exposure augmentation from tab. 2. We also randomly apply a tone-mapping operator between gamma, Filmic, or Hybrid-Log-Gamma to ensure the model is robust to various camera functions. We normalize the output of the tonemapping operators to [-$[\text{-}1,1]$ to fit Stable Diffusion’s original training range. We obtain the prompts from each image using LLaVA [35].
Parameter Distribution
Azimuth $\mathcal{U}[0,2\pi]$
Elevation $\mathcal{U}[-\nicefrac{{\pi}}{{2}},\nicefrac{{\pi}}{{2}}]$
Roll $\mathcal{U}[-\nicefrac{{\pi}}{{16}},\nicefrac{{\pi}}{{16}}]$
Vert. FoV $\mathcal{U}[60^{\circ},120^{\circ}]$
Exposure $2^{e},e\sim\mathcal{U}[-8,4]$
Table 2: Sampling parameters to generate training data for our HDR light intensity estimation network. Here, $\mathcal{U}[a,b]$ denotes a uniform distribution in the $[a,b]$ interval.
Our captured data was acquired in static scenes using a Ricoh Theta Z1 with exposure bracketing of 10 captures spanning from 1s to 1/25000s to obtain unsaturated captures. The camera was set to $f/2.1$, ISO 80, and fixed 5000K white balance for all captures. We stitched the RAW stereo spherical captures using PTGui [42].
We initialize our Stable Diffusion 1.5 backbone [50] with the IC-Light [73] checkpoint and fine-tune it for 415k iterations. We employ an initial learning rate of $10^{\text{-}5}$. Training took 10 days on 8 A100 GPUs with 40GB of VRAM. We perform 20 diffusion steps at inference, which takes a total of approximately 40 seconds on an Nvidia Titan V GPU.
3.2 Gaussian Splat lighting representation
Figure 2: Overview of our 3D reconstruction pipeline. From a collection of LDR input images, we first convert them to HDR by using our HDR light intensity estimation network (bottom, see sec. 3.1). A 3D Gaussian Splatting (3DGS) representation is subsequently trained on the resulting HDR images (top).
We utilize the 3D Gaussian Splatting framework (3DGS) [29] as our HDR lighting representation. This representation offers multiple benefits, establishing it as a superior option over other lighting representations, as detailed in tab. 1.
As illustrated in fig. 2, we first process each image in the set with our HDR extrapolation network (c.f. sec. 3.1) to obtain (linear) HDR images. Then, we apply the training process proposed by InstantSplat [14] to recover an HDR 3DGS representation of the scene (in linear space). To train the 3DGS, the following loss is used:
$\ell=\ell_{2}(\log(y^{*}+1),\log(y+1))\,,$ (1)
where $y$ and $y^{*}$ are the predicted and ground truth HDR pixel values, respectively.
4 Evaluation
In this section, we first discuss the datasets used for evaluation, followed by an evaluation of our light source intensity estimation model against previous LDR-to-HDR methods. Lastly, we demonstrate that our 3D lighting framework better represents lighting conditions in natural scenes than previous work.
4.1 Data
Refer to caption Refer to caption
Refer to caption Refer to caption
(a) Si-HDR BtP-HDR
Figure 3: Comparison of images from the (a) Si-HDR [23] and (b) BtP-HDR [5] datasets. Image crops are shown on the top row, while the bottom row shows the corresponding plot of the luminance, showing of (a) a saturated sample from the Si-HDR dataset; and (b) an unsaturated sample from BtP-HDR. The large flat regions of maximum intenstiy in (a) correspond to saturated pixel values. the BtP-HDR dataset does not contain such saturation.
To evaluate our light source intensity estimation model from sec. 3.1, we employ the following two datasets from the literature.
Si-HDR [23]
is comprised of 195 captures, each composed of 5 RAW exposures, which are merged into HDR stacks. They also provide a reference LDR image for each capture, which is obtained by applying a custom CRF and adding realistic camera noise modeled from the Canon 5D Mark III. The dataset depicts various indoor and outdoor scenes with light sources and saturated highlights in the frame.
Analyzing the dataset reveals that 82 captures hold saturated pixels, even at the fastest shutter speed, an example of this is shown in fig. 3. While these captures allow for expanding the dynamic range of LDR images for purposes such as re-tonemapping, the saturated captures cannot faithfully represent the dynamic range of light sources. In our evaluation, we eliminate the saturated captures, which results in 113 HDR images.
Beyond the Pixel [5] (BtP-HDR)
The unsaturated HDR dataset proposed for evaluation in [5] was expressly conceived to evaluate the presence of bright light sources in images. Contrarily to Si-HDR, the LDR reference image of this dataset is an actual capture from a camera, instead of a simulated re-exposition from the HDR merged image. We dub this dataset BtP-HDR henceforth.
4.2 Light source estimation
Si-HDR BtP-HDR
Method MSE PSNR MSE PSNR
LDR $0.0715$ $13.0605$ $0.0282$ $17.5467$
HDRCNN $0.0246$ $23.6138$ $0.0060$ $30.6715$
MaskHDR $0.0243$ $23.5503$ $0.0061$ $30.9510$
SingleHDR $0.0178$ $24.0680$ $0.0092$ $23.6101$
ExpandNet $0.0187$ $24.7158$ $0.0148$ $22.5799$
Exposure Diff $0.0264$ $23.7384$ $0.0061$ $27.4963$
Ours $0.0131$ $25.3895$ $0.0041$ $28.7022$
Table 3: Quantitative evaluation of renderings lit by each method’s HDR image output. Results are color coded by best, second and third best.
We evaluate our novel light intensity estimation method (sec. 3.1) on the task of bright light source prediction. We first show qualitative results on the filtered Si-HDR dataset in fig. 4 and on the BtP-HDR dataset in fig. 5. The figures illustrate both the luminance of the generated HDR maps, as well as a render of a sphere on a plane lit by the color HDR map. Since these images are not full 360° environment maps, we project them onto a partial hemisphere of 90° field of view, and use this as light source for rendering. The camera, plane, and sphere are located at the origin, facing the camera. The glossy sphere has an albedo of 1 (pure white) and a roughness coefficient of 0. We make sure all HDR images used as lighting are exposed the same way by aligning the averages of pixels with intensity between $0.2$ and $0.8$ with the ground truth. Observe how the predictions made by our method more closely match the ground truth, especially when bright light sources generate hard cast shadows on the plane and bright specularities on the sphere.
To further assess the accuracy of our method, we compute quantitative image comparison metrics (MSE and PSNR) on the renders compared to the ground truth in tab. 3. We observe that our method significantly outperforms existing methods in terms of both MSE and PSNR on the filtered Si-HDR dataset. On the BtP-HDR dataset, our method provides the best MSE score while offering a competitive PSNR value. We attribute this lower score to the fact that BtP-HDR features indoor scenes that tend to have a lower dynamic range than outdoor environments. Consequently, techniques like HDRCNN and MaskHDR perform well in areas that are slightly above saturation.
4.3 3D lighting representation
PanoHDR-NeRF Ours
Scene MSE PSNR MSE PSNR
Chess room $0.0247$ $17.5010$ $0.0053$ $25.9313$
Small class $0.1097$ $9.7872$ $0.0061$ $23.1244$
Table 4: Quantitative evaluation of renderings lit by novel views of PanoHDR-NeRF [19] and Ours, as compared to the captured HDR ground truth panoramas provided in [19]. Our method achieves the best scores on both scenes for all metrics.
Input LDR HDRCNN MaskHDR SingleHDR ExpandNet Exposure Diff ours GT
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Qualitative evaluation on the filtered Si-HDR dataset [23]. From an LDR image (left), we compare the luminance of each method’s predicted HDR (even rows) and the rendering of a glossy sphere illuminated by the prediction (odd rows). Our approach more accurately estimates light intensity, producing renders similar to those rendered using the ground truth HDR image.
Input LDR HDRCNN MaskHDR SingleHDR ExpandNet Exposure Diff ours GT
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Qualitative evaluation on the BtP-HDR dataset [5]. From an LDR image (left), we compare the luminance of each method’s predicted HDR (even rows) and the rendering of a glossy sphere illuminated by the prediction (odd rows). Our approach more accurately estimates light intensity, producing renders similar to those rendered using the ground truth HDR image.
The approach that is most closely related to ours is that of PanoHDR-NeRF [19], who train a NeRF-based implicit representation on images that have been converted to HDR using the LANet [69] LDR-to-HDR architecture. They also provide a dataset of interior scenes captured with a 360° camera, in which a few (between 8–10) ground truth HDR panoramas were captured using exposure bracketing.
We therefore compare our HDR Gaussian Splatting representation to PanoHDR-NeRF on two scenes from their dataset and show the quantitative results in tab. 4. These metrics were computed by first rendering a synthetic scene (similar to sec. 4.2), lit by each method’s prediction, and compared to the ground truth. The table shows that our method significantly outperforms the baseline. We note that the results they report in their paper also include a version of LANet fine-tuned on a small dataset of HDR images captured by the same camera that was used to capture the ground truth—we use the “vanilla” model to avoid unfairly advantaging the method, and assume that no method has knowledge of the actual camera used to capture the data.
5 Application: virtual object insertion
The 3DGS representation is inherently 3D, allowing for near-field lighting effects out-of-the-box. It is possible to composite virtual objects directly in this representation, as shown in LABEL:fig:teaser. We obtain these results by importing our 3DGS representation directly in Blender [8], setting the Gaussians as emitters, and rendering the scene using the Cycles physically-based renderer. The inserted objects are automatically lit with near-field lighting effects, without manual intervention or custom lights in the scene.
6 Conclusion
Limitations
Despite encouraging results, GaSLight suffers from some limitations. First, each Gaussian in the 3DGS representation is effectively considered to be a light source (e.g., when rendering, see sec. 5. This precludes the simulation of more complex lighting effects, such as interreflections or scattering. Here, it is likely that new raytracing-based approaches [40] could help. Second, our HDR light intensity estimation network is not trained to achieve multi-view consistency: it is therefore possible that different content is extrapolated for two nearby views. While the spherical harmonics in the 3DGS representation alleviates for this somewhat and can realistically blend inconsistent predictions, future work could enforce such consistency, for example through score distillation sampling [46]. We also note that, as opposed to most lighting estimation methods that predict light sources outside the field of view of the image, we only recover the lighting for visible parts of the scene. Combining our method with view extrapolation approaches (e.g., [7, 43]) is a promising future direction. Finally, our current implementation, which relies on InstantSplat [14] does not scale well to large scenes. This is somewhat orthogonal to our goal of capturing lighting—we are hopeful that new developments (e.g., [30]) could help alleviate this.
We present GaSLight, a novel Gaussian Splatting-based HDR Lighting representation. The core of our method is a diffusion-based LDR-to-HDR model that uses its priors of natural images to expand the dynamic ranges of LDR images to unsaturated, even on bright light sources such as the sun. We further fit a 3DGS model on the predicted HDR images, which results in a light representation that can model spatially-varying effects, achieves high frequency reflections and hard cast shadows, and can readily be integrated in existing rendering engines.
Acknowledgements  This research was supported by Sentinel North, NSERC grant RGPIN 2020-04799, and the Digital Research Alliance Canada.